XYZ Company uses a job order cost system to determine the cost of doing repair work. The average shop direct labor rate is $20 per hour. Estimated costs and expenses for the coming year are as follows: Engine Parts $675,000 Shop direct labor 300,000 Depreciation on shop equipment 65,000 Shop supervisor salaries 115,000 Shop property taxes 33,000 Shop supplies 12,000 Advertising expense 40,000 Administrative office salaries 87,000 Administrative office depreciation expense 10,000 Instructions; Calculate the predetermined overhead rate using direct labor hours as the activity base.
Answer #1

Direct labor hours = 300,000 /20 = 15000 hours

Depreciation on shop equipment 65,000
Shop supervisor salaries 115,000
Shop property taxes 33,000
Shop supplies   12,000
Total overheads 225000

Pre determined overhead rate = 225,000 / 15000 = $ 15 per direct labor hour